Green Bank, WV
April 16,,  2002

TO: J. Acree, C. Beaudet, R. Fisher, J. Ford, F. Ghigo, R. Lacasse,  R. McCullough, C. Niday, W. Sizemore
FROM: R. Fisher
SUBJ.: Minutes of Interference Protection Group 4-16-02 Meeting

In attendance: Acree, Beaudet, Fisher, McCullough, Niday, Sizemore; later by Ford and Lacasse

Minutes of Previous Meeting

 New action list and agenda format:
Jeff presented a new format for tracking action items and documenting progress on each item. The list will be maintained in PDF format on the web and and made available upon request in Word Perfect format and hard copy. Most action items currently listed in the monthly meeting agenda will be transferred to the action item list, thus bringing the agenda back to a manageable size, providing better tracking of each item, and reducing the burden of meeting minutes write-ups. Major engineering projects, such as the GBT feed arm servo RFI suppression and shield room investigation, will continue as agenda items since they are too extensive for action item documentation. Also, short term actions of just a few days duration will generally not be tracked on the action item list.
The columns of the new action item list are: item number, short name, description, proposed action, progress, point-of-contact (POC), created and due dates, and status (open or closed). Closed items will become invisible but will remain accessible for recall as needed. The question of how action item priorities will be handled was raised but not resolved. The POCs will be responsible for supplying a brief progress report in written and oral form at the time of the monthly IPG meeting.
 Status of current action items:
These minutes will document the action item discussions at this meeting as a transition to the new scheme. At the next meeting only topics specifically called out on the agenda will appear in the minutes, except by reference. Please see the current action item list under "Current action items" on the "RFI Protection" web page under "Resources" on the GB Home Page for the status of open items. The action item numbers below refer to the number in the new list, not to the order in which they appear in this meeting's agenda.

        1 - RFI Database (Acree, Ford)

Jeff is maintaining a spreadsheet for high-risk RFI. John is working on an automated database.
        2 - Portable RFI Sniffing Setup (Acree, Beaudet)
The following have been secured: Portable antennas, tripods, amplifiers, RF cables, adapter kit, and a high performance portable spectrum analyzer. The portable setup provides continuous coverage from DC - 18 GHz. Miscellaneous needs such as a long extension cord, cart, 250 ft. rule and laser distance finder have also been moved to the chamber area. We still need a dedicated low-cost plotter or printer.
        3 - System Controller PC at 40" (Beaudet)
A few modifications were made the the PC used at the 40-ft telescope, and it was returned to service. A reminder to turn off this computer when not in use may be in order. There are no other current plans for action on this item unless a Tempest enclosure can be procured.
        4 - MIT Research PC (Beaudet)
Item closed.
        5 - Fibercom Whisperlan Transceivers (Beaudet)
A prototype shielded box design for the replacement Centrecoms has been completed. Fabrication underway. Installation pending. Ed is planning to install the Centrecoms this summer. We may want to wait for the completion of the shielded boxes. The required parts are on order.
        6 - Turntable for Chamber (Beaudet)
Due to design and schedule challenges, a second positioner, independent of the indoor antenna range will be pursued. A load-bearing turntable, for use with the existing SA positioner, will be built in house. The SA positioner and controller have been moved to the chamber area and checked out ok. Carla is working on the mechanical design and floor plan. This action will probably be too big to continue to track on the AI list. The drawings are in Rob Taggart's hands, and the details of the concrete base and its compatibility with the antenna range turret are under study.
        7 - Atmospheric Phase Interferometer (API) (Beaudet)
Tests of DC and AC filtering were successful, but the enclosure may undergo a redesign due to radiation from the ribbon cable. Galen will make a decision on redesign at a date to be specified.
        8 - Antenna Roundup (Sizemore)
Wes sent an all hands email in to get things started. The equipment found falls into two categories: personal equipment (FM receivers, scanners, walky-talkies, etc.) and official use (GPS timing receiver, unicom, site two-way, API, weather radio, deer stand, lab FM radio distribution cable, etc.). The lab FM radio distribution cable is in need of major rework. Thoughts on policy for personal equipment was requested within the next few days.
        9 - Weather Radio (Acree)
A bandpass filter with 3dB points at 150 MHz and 170 MHz was installed at the antenna input. The reception of the 162.455 MHz weather frequency was not affected. The second harmonic, (151.115 MHz) was not suppressed and continues to be a potential RFI risk. It will be added to the RFI spreadsheet and if it shows up on a GBT RFI report, additional steps will be taken. This item is now closed.
        10 - HVAC 43 MHz RFI (Beaudet, Niday)
Emissions were seen at frequencies other than 43 MHz from this unit, although 43 MHz is the most prominent. Bob Goldizen will be contacted to begin working out a strategy for shielding and filtering. Another copy of the unit would be very useful for measurements and suppression tests.
        11 - 370 MHz wideband noise toward Arbovale (Acree)
Nothing was found in a survey with the interference truck. Rick will look again with the deer stand antenna to see if the emission is still present.
        12 - Boltec Lightning Detector (Niday)
Chuck finally got a response from the manufacturer, who said that all signals on the four-pair cable to the antenna unit were below one MHz. Design of a low-pass filter set is under way.
        13 - FM Signal Distribution System & Antenna (Niday)
Several tests were done. Found a source of bandpass filters. The whole system needs upgrade.
        14 - DGS Laser System data acquisition module
This device was tested in the anechoic chamber, and significant RFI was measured. Chuck will supply data on these measurements. The fix was left to the metrology group, but Randy will follow up on the current status as part of his laser RFI suppression efforts.
        15 - Blue Warning Lamps (Ford)
Requires documentation in action list.
        16 - Storage Area (Sizemore)
After much searching and negotiation, Wes secured the storage shed that was previously used for storing the subreflector for the 140 ft. radio telescope. The hardware has been move and organized for cataloging and allocation. There is a lot of useful hardware in this collection, and some decisions on what to keep in the IPG inventory and what to turn into the general electronics stock are required. Within a few months a new storage area is required since the current area is needed for oil storage.
        17 - GlobalStar, Iridium, Boeing - Connexion (Acree)
Jeff has the documents that explain the process and has discussed the matter with Mark McKinnon. Some discussion ensued about how coordination might fit into dynamic scheduling and the need for publishing schedules to the satellite operators to maintain visibility.
        18 - Notebook Computers and Network Connections
The converter box and a notebook PC need to be tested in the anechoic chamber. Need to determine whether radiation from the converter box is significantly stronger that the notebook emissions. If not, one 'turn off when not in use' policy may suffice. Otherwise, the boxes may require suppression.
        new - Auditorium projection rack (Beaudet)
A cover plate and switches are on order for this system. After installed the system will be labeled with a 'Turn me off when not in use' sign.
        new - PBX ground (Beaudet)
No significant RFI suppression was achieved with a trial water pipe grounding of this rack. We generally agreed that a three-ground-rod system with a common point as close to the PBX ground as possible is a desirable goal. Carla will continue to work with the plant maintenance people.
        new - LAN equipment shielding (Ford)
Ed Childers is making progress on getting the two LAN systems into RFI-tight enclosures. Some of the downstairs system is being eliminated and the rest moved into a shielded box. The upstairs system needs a shielded rack.
        new - Deer stand emission check (Fisher)
No progress on determining whether any RFI seen at the deer stand could be generated by the deer stand equipment itself. Rick set a completion date of May 30.
        new - Tour center fluorescent lights (Beaudet)
The fluorescent lights to be installed in the new tour center have been provided by the contractor for anechoic chamber measurement. This and similar devices for approval will become action items only if they prove problematic. The need to document tests with specific model numbers was noted as a control on changes to the contractor's installation work. Copies of test results will be supplied to Butch Wirt, who is overseeing the contract.
        new - Blue warning lamps (Ford)
Not discussed
 Major projects and other agenda items:

        GBT feed arm servo (McCullough, Lacasse)

The initial shipment of filters for tests is due arrive in two weeks. Two additional weeks will then be required for the tests. The big box that will house the filters failed its suppression test and has gone back to the shop for the installation of more fasteners. The shielded enclosures for the magnetorestrictive transducers are ready for installation, possibly during the shutdown for track grout rework at the beginning of May.
        Lasers (McCullough)
The conductive paint adhesion problem was traced to the use of a paint stripper that was incompletely removed. The enclosure was thoroughly cleaned with acetone and all remaining old paint removed mechanically. The new conductive paint application was completely successful. All required parts except a flat gasket are in house ready for assembly of the larger RFI enclosure. Test on leakage of 1500 MHz radiation from the smaller enclosure found that the signal cables had been adequately filtered in the original design and that a only mechanical improvement of the shielding is required.
        GBT main drive emissions (Acree, Beaudet, Ford, Fisher)
Measurements of RFI from the main GBT drives with an antenna stationed at the warehouse near the GBT found only modest radiation from the power switching circuits. Nothing consistent with the report of strong RFI at the 40-ft telescope was found. Subsequent tests at the 40-ft showed no significant RFI when the GBT was slewing, as had originally been reported. It is possible that efforts to improve the filtering of cables passing through the main GBT servo room walls suppressed the RFI seen at the 40-ft. This item is dormant unless further cause for concern emerges.
        Lab shielded room integrity (Norrod, Ford)
Extensive measurements and materials investigations were reported by Roger in an email of Thursday, 11 Apr 2002. A brief summary is that the shielding effectiveness of these rooms has degraded from the original value of >60 dB to the range of 30 to 40 dB due to corrosion around the window flanges and apparent chemical action which has eaten away much of the wall paper copper. The window flanges are being reworked, and further measurements of the degraded wall paper are under way. Carla reported in emails of Thursday 11 Apr 2002 15:02:31 and Monday 15 Apr 2002 08:28:19 roughly 30 to 50% correspondence between RFI signal frequencies measured inside the shielded room and signal frequencies measured with the deer stand and the GBT in the 300-400 MHz range. Very little coincidence was found with RFI signal frequencies measured in the 775-825MHz range on the GBT
 Next Meeting:  5 - 21 - 02, 0800